# Fast Diffusion-Based Counterfactuals for Shortcut Removal and Generation

### v1

@@ Appendix D Denoised images for efficient gradient estimation
Appendix E Mask-based inpanting
Initially, denoised images used for gradient estimation are blurred. We found both quantitative (see Appendix C) and qualitatively that we need to start applying our mask-based inpainting strategy as long as denoised images are not blurred, i.e, at $\tau_{w}=\frac{\tau}{2}$. An example of how our method FastDiME progresses across different time steps is illustrated in Fig. 15, together with the thresholded masks $M_{t}$ used during our inpainting scheme and the final difference map between the counterfactual and the original image. To highlight its importance, we also provide visual examples of our method without the mask-based inpainting (FastDiME (w/o Mask)) in Fig. 16 together with some failure cases of our method.
@@ Appendix E Mask-based inpanting
Figure 8: CelebA. Comparisons for the ‘smile’ attribute
@@
Figure 9: CelebA. Comparisons for the ‘age’ attribute
@@
Figure 15: Maks-based inpainting. An example of how our method FastDiME progresses across different time steps. We show the path toward the final counterfactual for two examples (left: no smile normal-$\rightarrow$ smile, right: smile normal-$\rightarrow$ no smile). We start applying our mask-based inpainting strategy from $\tau_{w}=30$. To better highlight the differences between steps we show the threshold masks before applying dilation.
@@
Figure 16: CelebA. Examples for no smile normal-$\rightarrow$ smile counterfactuals. We show examples of our method without the mask-based inpainting (FastDiME (w/o Mask)). In most cases, our mask-based inpainting strategy helps to improve the results (top). However, there are still some hard examples where it is not always successful (bottom).

### v2

@@ Appendix 0.D Denoised images for efficient gradient estimation
Appendix 0.E Self-optimized masking
Initially, denoised images used for gradient estimation are blurred. We found both quantitatively (see Appendix 0.C) and qualitatively (see Appendix 0.D) that we need to start applying our self-optimized masking scheme as long as denoised images are not blurred, i.e, at $\tau_{w}=\frac{\tau}{2}$. An example of how our method FastDiME progresses across different time steps is illustrated in Fig. 15, together with the thresholded masks $M_{t}$ used during our self-optimized masking scheme and the final absolute difference map between the counterfactual and the original image. To highlight its importance, we also provide visual examples of our method without the self-optimized masking scheme (FastDiME (w/o Mask)) in Fig. 16.
@@ Appendix 0.E Self-optimized masking
Figure 8: CelebA. Comparisons for the ‘smile’ attribute.
@@
Figure 9: CelebA. Comparisons for the ‘age’ attribute.
@@
Figure 15: Self-optimized masking. An example of how our method FastDiME progresses across different time steps. We show the path toward the final counterfactual for two examples (left: no smile $\rightarrow$ smile, right: smile $\rightarrow$ no smile). We start applying our self-optimized masking scheme from $\tau_{w}=30$. To better highlight the differences between steps we show the thresholded masks before applying dilation.
@@
Figure 16: CelebA. Examples for no smile $\rightarrow$ smile counterfactuals. We show examples of our method without the self-optimized masking scheme (FastDiME (w/o Mask)). Our self-optimized masking scheme helps to improve the results.